Con you moke devotion step step)? Fomutos. ( Please do not wlite only step by step. Because no formulas Jou can fnd theee formblasExplan th oree as the cmas vot chae testic Pleo ose exp lah ceep
Answer #1

Consider a CMOS Inveter circuit i,e Combination of both PMOS and NMOS

Assume NMOS is On(Saturated),PMOS is linear

Then Drain current expressed as follows

DS tVDS

I_{d}=K\left [ (V_{GS}-V_{t})V_{DS}-\frac{V_{DS}^{2}}{2} \right ]

which can be modified as follows for above assumed conditions for both N and P MOS as follows

I_{d}=K_{p}\left [ \left ( V_{dd}-V_{i} -V_{tp} \right )\left ( V_{dd}-V_{o} \right )-\frac{1}{2}\left ( V_{dd}-V_{o} \right )^{2}\right ]=\frac{K_{n}}{2}\left ( V_{i}-V_{tn} \right )^{2}

let

V_{dp}=\left ( V_{dd}-V_{o} \right )

\beta =\frac{K_{n}}{K_{p}}

dp2K

\Rightarrow \left [ \left ( V_{dd}-V_{i} -V_{tp} \right )V_{dp} -\frac{1}{2}V_{dp}^{2}\right ]=\frac{\beta }{2}\left ( V_{i}-V_{tn} \right )^{2}

\Rightarrow \frac{1}{2}V_{dp}^{2}-\left ( V_{dd}-V_{i} -V_{tp} \right )V_{dp} +\frac{\beta }{2}\left ( V_{i}-V_{tn} \right )^{2}=0

Obtaining solution for above equation

\alpha _{1},\alpha _{2}=\frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

\alpha _{1},\alpha _{2}=\frac{\left ( V_{dd}-V_{i} -V_{tp} \right )\pm \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}-4*\frac{1}{2}*\frac{\beta }{2}\left ( V_{i}-V_{tn} \right )^{2}}}{2*\frac{1}{2}}

\alpha _{1},\alpha _{2}=\left ( V_{dd}-V_{i} -V_{tp} \right )\pm \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}-\beta \left ( V_{i}-V_{tn} \right )^{2}}

ddVi -

V_{dd}-V_{0}=\left ( V_{dd}-V_{i} -V_{tp} \right )- \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}-\beta \left ( V_{i}-V_{tn} \right )^{2}}

\Rightarrow V_{0}=V_{dd}-\left ( V_{dd}-V_{i} -V_{tp} \right )+\sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}-\beta \left ( V_{i}-V_{tn} \right )^{2}}

\Rightarrow V_{0}=V_{i} +V_{tp} + \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}-\beta \left ( V_{i}-V_{tn} \right )^{2}}

If K_{n}=K_{p}=K\Rightarrow \beta =1

\Rightarrow V_{0}=V_{i} +V_{tp} + \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}- \left ( V_{i}-V_{tn} \right )^{2}}

Finding VOH and ViL

\Rightarrow V_{0}=V_{i} +V_{tp} + \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}- \left ( V_{i}-V_{tn} \right )^{2}}

\Rightarrow V_{0}=V_{i} +V_{tp} + \sqrt{\left ( V_{dd}-V_{i} -V_{tp}+V_{i}-V_{tn}\right )\left ( V_{dd}-V_{i} -V_{tp}- \left ( V_{i}-V_{tn} \right )\right )}

\Rightarrow V_{0}=V_{i} +V_{tp} + \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )}\rightarrow 1

we Know that

\frac{\partial V_{0}}{\partial V_{in}}=-1

\frac{\partial\left (V_{i} +V_{tp} + \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )} \right )}{\partial V_{in}}=-1

1+0+

\Rightarrow \frac{ 1 }{2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )}}*\left ( V_{dd} -V_{tp}-V_{tn}\right )(-2)=-2

\Rightarrow \frac{ 1 }{2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )}}*\left ( V_{dd} -V_{tp}-V_{tn}\right )=1

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )=2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )}

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )^{2}=4\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )=4\left ( V_{dd}-2V_{i} -V_{tp}+V_{tn} \right )

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )=\left (4 V_{dd}-8V_{i} -4V_{tp}+4V_{tn} \right )

\Rightarrow 8V_{i}=\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )

\Rightarrow V_{iL}=\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}\rightarrow 2

Substitute eqn 2 in 1 we will get

\Rightarrow V_{0H}=\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}+V_{tp} + \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-2(\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}) -V_{tp}+V_{tn} \right )}\Rightarrow V_{0H}=\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}+V_{tp} + \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( V_{dd}-\frac{3 V_{dd} -3V_{tp}+5V_{tn} }{4} -V_{tp}+V_{tn} \right )}\Rightarrow V_{0H}=\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}+V_{tp} + \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( \frac{4V_{dd}-3 V_{dd} +3V_{tp}-5V_{tn}-4V_{tp}+4V_{tn} }{4} \right )}

\Rightarrow V_{0H}=\frac{\left (3 V_{dd} -3V_{tp}+5V_{tn} \right )}{8}+V_{tp} + V_{dd} -V_{tp}-V_{tn}

Finding VOL and ViH

\Rightarrow V_{0}=V_{i} -V_{tn}- \sqrt{\left ( V_{dd}-V_{i} -V_{tp} \right )^{2}- \left ( V_{i}-V_{tn} \right )^{2}}

\Rightarrow V_{0}=V_{i} -V_{tn}- \sqrt{\left ( V_{dd}-V_{i} -V_{tp}+V_{i}-V_{tn}\right )\left ( V_{dd}-V_{i} -V_{tp}- \left ( V_{i}-V_{tn} \right )\right )}

\Rightarrow V_{0}=V_{i} -V_{tn} -\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )}\rightarrow 3

we Know that

\frac{\partial V_{0}}{\partial V_{in}}=-1

\frac{\partial\left (V_{i} -V_{tn} -\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )} \right )}{\partial V_{in}}=-1

1-0-\frac{\partial\left ( \sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )} \right )}{\partial V_{in}}=-1

\Rightarrow \frac{ -1 }{2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )}}*\left ( V_{dd} -V_{tp}-V_{tn}\right )(2)=-2

\Rightarrow \frac{ -1 }{2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )}}*\left ( V_{dd} -V_{tp}-V_{tn}\right )=-1

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )= 2\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )}

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )^{2}= 4\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )=4\left ( -V_{dd}+2V_{i} +V_{tp}-V_{tn} \right )

\Rightarrow \left ( V_{dd} -V_{tp}-V_{tn}\right )=\left ( -4V_{dd}+8V_{i} +4V_{tp}-4V_{tn} \right )

\Rightarrow \left ( 5V_{dd} -5V_{tp}+3V_{tn}\right )=8V_{iH}

\Rightarrow V_{iH}=\frac{\left ( 5V_{dd} -5V_{tp}+3V_{tn}\right )}{8} \rightarrow 4

Substitute equation4 in equation3

\Rightarrow V_{0L}=\frac{\left ( 5V_{dd} -5V_{tp}+3V_{tn}\right )}{8} -V_{tn} -\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right )\left ( -V_{dd}+2\left ( \frac{\left ( 5V_{dd} -5V_{tp}+3V_{tn}\right )}{8} \right )+V_{tp}-V_{tn} \right )}

\Rightarrow V_{0L}=\frac{\left ( 5V_{dd} -5V_{tp}+3V_{tn}\right )}{8} -V_{tn} -\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right ) \left ( \frac{ -4V_{dd}+5V_{dd} -5V_{tp}+3V_{tn}+4V_{tp}-4V_{tn}}{4} \right ) }

\Rightarrow V_{0L}=\frac{\left ( 5V_{dd} -5V_{tp}-5V_{tn}\right )}{8} -\sqrt{\left ( V_{dd} -V_{tp}-V_{tn}\right ) \left ( \frac{ -4V_{dd}+5V_{dd} -5V_{tp}+3V_{tn}+4V_{tp}-4V_{tn}}{4} \right ) }\Rightarrow V_{0L}=\frac{\left ( 5V_{dd} -5V_{tp}-5V_{tn}\right )}{8} -\left ( V_{dd} -V_{tp}-V_{tn}\right )

\Rightarrow V_{0L}=\frac{\left ( 5V_{dd} -5V_{tp}-5V_{tn}\right )-8\left ( V_{dd} -V_{tp}-V_{tn}\right )}{8}

VOL